# Cosmology in Modified $f(\mathcal{G})$ Gravity: A Late Time Cosmic Phenomena

### v1

@@ III.2 Supernovae type Ia (SNe Ia)
We assess the models against the standard $\Lambda$CDM model using the Akaike Information Criterion (AIC) [44] and the Bayesian Information Criterion (BIC) [45] in addition to $\chi^{2}_{\text{min}}$. Both AIC and BIC take into account the model’s goodness of fit as well as its complexity, which is influenced by the number of parameters $(n)$. The AIC is determined as
$AIC=\chi^{2}_{\text{min}}+2n$ (19)
@@
$BIC=\chi^{2}_{\text{min}}+n\,\text{ln}\,\mathcal{N}$ (20)
@@
$\text{AIC}_{\text{c}}=\text{AIC}+\frac{2n(n+1)}{\mathcal{N}-n-1}$ (21)
@@
Figure 1: Contour plots show the $1-\sigma$ and $2-\sigma$ uncertainty areas for the variables $H_{0}$, $\alpha$, $\beta$, and $m$. These contours are derived from the CC sample (left panel) and the Pantheon+ data (right panel).
Figure 2: In the upper panel, the black error bars show uncertainty for 32 data points from the CC sample, with the solid blue line representing the model and the broken red line representing $\Lambda$CDM. In the lower panel, the solid blue line represents the model’s distance modulus $\mu(z)$ against redshift $z$, providing a superior fit to the 1701 data points from the Pantheon+ dataset with error bars.
Datasets $H_{0}$ $\Omega_{m0}$ $\alpha$ $\beta$ $m$
CC Sample $68.944^{+1.210}_{-1.551}$ $0.355^{+0.043}_{-0.014}$ $464.593^{+464.593}_{-415.110}$ $47573.273^{+43434.714}_{-42187.475}$ $10.342^{+7.984}_{-8.144}$
Pantheon+ $72.269^{+0.199}_{-0.197}$ $0.419^{+0.018}_{-0.019}$ $502.193^{+339.927}_{-340.285}$ $49572.389^{+34109.579}_{-33769.153}$ $10.412^{+6.488}_{-6.509}$
Table 1: The table above presents an exploration of the parameters for the MCMC algorithm. It displays the best-fit values for the model parameters, including $H_{0}$, $\alpha$, $\beta$, and $m$, derived from the MCMC study using the CC and Pantheon+ data sets.
@@
The contour plots (see Fig. 1) display the 1–$\sigma$ and 2–$\sigma$ uncertainty regions for the parameters $H_{0}$, $\Omega_{m0}$, $\alpha$, $\beta$, and $m$ using two datasets: the CC data and the Pantheon+ data. These plots, derived using the MCMC method, illustrate the marginalized posterior distributions of parameter pairs, with inner and outer contours representing $68\%$ and $95\%$ confidence levels, respectively.
@@
The contour plot for the parameters $H_{0}$ and $\Omega_{m0}$ exhibits a more elliptical shape compared to the other parameter pairs, which display more square or elongated contours. This elliptical shape indicates that there is a relatively weak correlation between $H_{0}$ and $\Omega_{m0}$, suggesting that the data independently constrains these two parameters. This independence implies that the variations in one parameter do not significantly affect the value of the other, leading to a more symmetrical uncertainty region. In contrast, the square or elongated contours seen in other parameter pairs indicate stronger correlations, where changes in one parameter can be offset by adjustments in another to maintain a similar fit to the data. This strong coupling results in less symmetrical and more stretched uncertainty regions, reflecting the interdependence of these parameters within the $f(\mathcal{G})$ gravity model. In order to evaluate the efficacy of our MCMC analysis, we calculated the associated AIC, $\text{AIC}_{\text{c}}$ and BIC values, which are presented in Table 2. Our findings strongly support the assumed $f(\mathcal{G})$ gravity models when analyzing the datasets. Moreover, we noted that the $f(\mathcal{G})$ model demonstrates greater precision when applied to the Pantheon+ datasets.
@@
The upper panel of Fig. 2 shows that the $f(\mathcal{G})$ gravity model fits the observational $H(z)$ data well across the redshift range considered. Both the $f(\mathcal{G})$ gravity and $\Lambda$CDM models follow similar trends, but the $f(\mathcal{G})$ model predicts lower $H(z)$ values at higher redshifts ($z>1$). This deviation suggests distinct underlying physics due to higher-order curvature terms in the $f(\mathcal{G})$ model, which also accounts for late-time cosmic acceleration without a cosmological constant ($\Lambda$).
The lower panel of Fig. 2 illustrates that the $f(\mathcal{G})$ gravity model fits the distance modulus $\mu(z)$ data excellently across the redshift range. Both the $f(\mathcal{G})$ gravity and $\Lambda$CDM models align closely with the observational data, with minimal deviation between them. The strong agreement with observational data supports its viability as a competitive alternative to the $\Lambda$CDM model, with its natural incorporation of higher-order curvature terms making it an attractive option for future cosmological studies.
Figure 3: Graphical representation of the deceleration parameter versus redshift using the constrained coefficients from Fig. 1.
III.3 Cosmological Parameter Evolution
@@
We analyze the evolution of crucial cosmological parameters, including the effective equation of state (EoS), statefinder, and Om diagnostic parameter, by imposing constraints on model parameters using various observational data.
Figure 3 presents the deceleration parameter $q$ as a function of redshift $z$ for the $f(\mathcal{G})$ gravity model, derived from both Hubble data (upper panel) and Pantheon+ data (lower panel). Fig. 5 demonstrate that the restricted values of model parameters derived from the analyzed CC and Pantheon+ data sets indicate a transition of $q$ from positive (indicating early deceleration) in the past to negative (indicating current acceleration) in the present. The present value of deceleration parameter $q_{0}$ is measured to be $-0.46$ and $-0.39$ for the CC and Pantheon+ data, respectively, at the current cosmic epoch, which aligns relatively well with the range of $q_{0}=-0.528^{+0.092}_{-0.088}$ determined by recent observations [46].
The consistency between the curves from Hubble and Pantheon+ data underscores the robustness of the $f(\mathcal{G})$ model in capturing the Universe’s expansion dynamics. This transition from near-zero values to negative values effectively illustrates how the model accounts for cosmic acceleration, driven by the higher-order curvature terms inherent in the $f(\mathcal{G})$ theory.
@@
Figure 4: Graphical representation of the EoS parameter versus redshift using the constrained coefficients from Fig. 1.
Figure 4 shows the effective equation of state parameter $w_{\text{eff}}$ as a function of redshift $z$ for the $f(G)$ gravity model. At low redshifts ($z\approx 0$), $w_{\text{eff}}$ is close to $-1$, indicating DE dominance and accelerated expansion. As $z$ increases, $w_{\text{eff}}$ transitions from negative values to less negative values, reflecting a shift from a deceleration phase in the early Universe to an acceleration-dominated phase in the current epoch. The asymptotic behavior of $w_{\text{eff}}$ at higher redshifts, approaching around $-0.3$, suggests significant early universe matter contribution, reducing DE’s influence over time.
At low redshifts ($z<0.5$), $w_{\text{eff}}$ starts near -1, consistent with the $\Lambda$CDM model, indicating that $f(\mathcal{G})$ gravity mimics a cosmological constant, leading to the observed accelerated expansion in the present. As the redshift increases ($1.50.5<z<$ 1.5), $w_{\text{eff}}$ rises, suggesting a transition influenced by $f(\mathcal{G})$ modifications. This indicates a slower acceleration rate compared to the standard model, emphasizing the distinct impact of $f(\mathcal{G})$ gravity during this period. At higher redshifts ($z>1.5$), $w_{\text{eff}}$ stabilizes around $-0.3$, indicating a reduced influence of $f(\mathcal{G})$ modifications in the early universe. This suggests a quintessence-like behavior, where the expansion is driven by a combination of matter and a less dominant DE component.
@@
The consistent behavior of $w_{\text{eff}}$ across both datasets reinforces the capability of the $f(\mathcal{G})$ model to describe the expansion history of the Universe. This transition aligns with theoretical expectations of the $f(\mathcal{G})$ model, which incorporates higher-order curvature terms to account for cosmic dynamics without a cosmological constant. The observed $w_{\text{eff}}$ behavior highlights the model’s effectiveness, supporting its viability as an alternative to the $\Lambda$CDM model.
@@
The statefinder diagnostic proposed by V. Sahni [47] provides a geometric method for discerning different DE models using statefinder parameters.
@@
Figure 5 illustrates the $r-s$ parameter plot for the $f(\mathcal{G})$ gravity model. The trajectory in the $r-s$ plane highlights the evolutionary track of the expansion of the Universe. The $f(\mathcal{G})$ model passes through the region corresponding to the $\Lambda$CDM model, indicated by the red point. At lower values of $s$, the model aligns with quintessence characteristics, suggesting a dynamical DE component with $\omega>-1$. As $s$ increases, the trajectory moves towards regions associated with Chaplygin gas models, indicating a unified dark matter and DE scenario.
The smooth transition observed in the $r-s$ parameter space demonstrates the flexibility of the $f(\mathcal{G})$ model in describing different cosmological behaviors. This capability allows the model to account for various dynamics, from quintessence-like to Chaplygin gas-like, providing a comprehensive description of the expansion of the Universe.
The Om diagnostic is a simple testing method that depends only on the first-order derivative of the cosmic scale factor. In particular, in DE theories, the $\text{Om}(z)$ parameter is followed as an additional effective diagnostic tool [48, 47] and is defined as,
$\text{Om}(z)=\frac{\frac{H^{2}(z)}{H^{2}_{0}}-1}{(1+z)^{3}-1}$ (23)
@@
$\text{Om}(z_{1},z_{2})=\text{Om}(z_{1})-\text{Om}(z_{2})$ (24)
Alternatively stated for simplification, when $\text{Om}(z_{1},z_{2})>0$, it indicates quintessence, whereas when $\text{Om}(z_{1},z_{2})<0$, it signifies phantom behavior, where $(z_{1}<z_{2})$. The $\text{Om}(z)$ diagnostic in the $\Lambda$CDM model serves as a null test, as noted in [48], and its sensitivity to the EoS parameter was further explored in subsequent data as seen in [49, 50, 51]. If $\text{Om}(z)$ remains constant for the redshift, the DE concept would be a cosmological constant. The DE concept will form a cosmological constant if $\text{Om}(z)$ is constant for the redshift. The slope of $\text{Om}(z)$, which is positive for the emerging $\text{Om}(z)$ and denotes phantom phase $(\omega<-1)$ and negative for quintessence region $(\omega>-1)$ also identifies the DE models.
@@
IV Stability analysis of dynamical systems
The techniques of dynamical systems are helpful in examining the overall long-term behavior of a specific cosmological model [52, 53, 54, 55, 56, 57] and can assist in circumventing the challenge of solving non-linear cosmological equations. These techniques characterize the universal dynamics by studying the nearby long-term behavior of critical points of the system and connecting them to the primary cosmological eras [58].
@@ IV Stability analysis of dynamical systems
$X=H^{2}f_{\mathcal{G}}\,,\quad Y=H\dot{f}_{\mathcal{G}}\,,\quad Z=\frac{\dot{H}}{H^{2}}\,,\quad$
$W=-\frac{f}{6H^{2}}\,,\quad V=\frac{\kappa^{2}\rho_{r}}{3H^{2}}\,.$ (25)
$\Omega_{\text{m}}=\frac{\kappa^{2}\rho_{m}}{3H^{2}},\quad\Omega_{\text{r}}=V=\frac{\kappa^{2}\rho_{r}}{3H^{2}},\quad\Omega_{\text{DE}}=\frac{\kappa^{2}\rho_{DE}}{3H^{2}},$ (26)

### v2

@@ 3.2 Supernovae type Ia (SNe Ia)
Figure 2: In the upper panel, the black error bars show uncertainty for 32 data points from the CC sample, with the solid teal line representing the model and the broken red line representing $\Lambda$CDM. In the lower panel, the solid teal line represents the distance modulus $\mu(z)$ of the model against redshift $z$, providing a superior fit to the 1701 data points from the Pantheon+ dataset with error bars.
We assess the models against the standard $\Lambda$CDM model using the Akaike Information Criterion (AIC) (Akaike, 1974) and the Bayesian Information Criterion (BIC) (Schwarz, 1978) in addition to $\chi^{2}_{\text{min}}$. Both AIC and BIC take into account the model’s goodness of fit as well as its complexity, which is influenced by the number of parameters $(n)$. The AIC is determined as
$\text{AIC}=\chi^{2}_{\text{min}}+2n\,,$ (22)
@@
$\text{BIC}=\chi^{2}_{\text{min}}+n\,\text{ln}\,\mathcal{N},$ (23)
@@
$\text{AIC}_{\text{c}}=\text{AIC}+\frac{2n(n+1)}{\mathcal{N}-n-1}\,,$ (24)
@@
Datasets $H_{0}$ $\Omega_{\text{m}0}$ $\alpha$ $\beta$ $m$
CC Sample $68.944^{+1.210}_{-1.551}$ $0.355^{+0.043}_{-0.014}$ $464.593^{+439.011}_{-415.110}$ $47573.273^{+43434.714}_{-42187.475}$ $10.342^{+7.984}_{-8.144}$
Pantheon+ $72.270^{+0.200}_{-0.201}$ $0.420^{+0.020}_{-0.020}$ $502.193^{+339.937}_{-340.285}$ $49572.390^{+34109.581}_{-33769.153}$ $10.412^{+6.491}_{-6.510}$
Table 1: The table above presents an exploration of the parameters for the MCMC algorithm. It displays the best-fit values for the model parameters, including $H_{0}$, $\Omega_{\text{m}0}$, $\alpha$, $\beta$, and $m$, derived from the MCMC study using the CC and Pantheon+ data sets.
@@
The contour plots (see Fig. 1) display the 1–$\sigma$ and 2–$\sigma$ uncertainty regions for the parameters $H_{0}$, $\Omega_{\text{m}0}$, $\alpha$, $\beta$, and $m$ using two datasets: the CC data and the Pantheon+ data. These plots, derived using the MCMC method, illustrate the marginalized posterior distributions of parameter pairs, with inner and outer contours representing $68\%$ and $95\%$ confidence levels, respectively.
@@
The contour plot for the parameters $H_{0}$ and $\Omega_{\text{m}0}$ exhibits a more elliptical shape compared to the other parameter pairs, which display more square or elongated contours. This elliptical shape indicates that there is a relatively weak correlation between $H_{0}$ and $\Omega_{\text{m}0}$, suggesting that the data independently constrains these two parameters. This independence implies that the variations in one parameter do not significantly affect the value of the other, leading to a more symmetrical uncertainty region. In contrast, the square or elongated contours seen in other parameter pairs indicate stronger correlations, where changes in one parameter can be offset by adjustments in another to maintain a similar fit to the data. This strong coupling results in less symmetrical and more stretched uncertainty regions, reflecting the interdependence of these parameters within the $f(\mathcal{G})$ gravity model. In order to evaluate the efficacy of our MCMC analysis, we calculated the associated AIC, $\text{AIC}_{\text{c}}$ and BIC values, which are presented in Table 2. Our findings strongly support the assumed $f(\mathcal{G})$ gravity models when analyzing the datasets. Moreover, we noted that the $f(\mathcal{G})$ model demonstrates greater precision when applied to the Pantheon+ datasets.
@@
The upper panel of Fig. 2 shows that the $f(\mathcal{G})$ gravity model fits the observational $H(z)$ data well across the redshift range considered. Both the $f(\mathcal{G})$ gravity and $\Lambda$CDM models follow similar trends, but the $f(\mathcal{G})$ model predicts lower $H(z)$ values at higher redshifts ($z>1$). This deviation suggests distinct underlying physics due to higher-order curvature terms in the $f(\mathcal{G})$ model, which also accounts for late-time cosmic acceleration without a cosmological constant ($\Lambda$). The lower panel of Fig. 2 illustrates that the $f(\mathcal{G})$ gravity model fits the distance modulus $\mu(z)$ data excellently across the redshift range. Both the $f(\mathcal{G})$ gravity and $\Lambda$CDM models align closely with the observational data, with minimal deviation between them. The strong agreement with observational data supports its viability as a competitive alternative to the $\Lambda$CDM model, with its natural incorporation of higher-order curvature terms making it an attractive option for future cosmological studies.
Figure 3: Whisker plot depicting the model parameters $H_{0}$, $\Omega_{\text{m}0}$, $\alpha$, $\beta$ and $m$, respectively with their $1-\sigma$ confidence intervals.
Our analysis reveals a significant discrepancy between the Hubble constant $H_{0}$ values derived from the CC sample and the Pantheon datasets, as shown in Fig. 3. This whisker plot highlights the ongoing $H_{0}$ tension in cosmology by presenting the model parameters $H_{0}$, $\Omega_{\text{m}0}$, $\alpha$, $\beta$, and $m$ along with their $1-\sigma$ confidence intervals.
3.3 Cosmological Parameter Evolution
@@
We analyze the evolution of crucial cosmological parameters, including the effective EoS, statefinder, and Om diagnostic parameters, by imposing constraints on model parameters using various observational data. We present a fully general expression for the deceleration parameter, $q=-\ddot{a}/aH^{2}$, as follows:
$q(z)=-1+\frac{H^{\prime}(z)}{H(z)}(1+z)\,.$ (25)
@@
Figure 4: Graphical representation of the deceleration parameter versus redshift using the constrained coefficients from Fig. 1. The thick line represents the behavior of the deceleration parameter for the $f(\mathcal{G})$ model, while the dashed line shows the deceleration parameter for the $\Lambda$CDM model.
Fig. 4 presents the deceleration parameter as a function of redshift $z$ for the $f(\mathcal{G})$ gravity model, derived from both Hubble data (upper panel) and Pantheon+ data (lower panel). Fig. 4 demonstrate that the restricted values of model parameters derived from the analyzed CC and Pantheon+ data sets indicate a transition of $q$ from positive (indicating early deceleration) in the past to negative (indicating current acceleration) in the present. The present value of deceleration parameter $q_{0}$ is measured to be $-0.527$ and $-0.499$ for the CC and Pantheon+ data, respectively, at the current cosmic epoch, which aligns relatively well with the range of $q_{0}=-0.528^{+0.092}_{-0.088}$ determined by recent observations (Gruber & Luongo, 2014). Recent observations are consistent with this deceleration parameter, and the resulting model indicates a smooth transition from deceleration to acceleration at $z_{t}=0.84$ and $z_{t}=0.82$ for the CC and Pantheon+ datasets, respectively. The derived transition redshift $z_{t}$ aligns with current constraints based on 11 $H(z)$ observations reported by Busca et al. (2013) for redshifts $0.2\leq z\leq 2.3$, $z_{t}=0.74\pm 0.5$ from Farooq & Ratra (2013), $z_{t}=0.7679^{+0.1831}_{-0.1829}$ by Capozziello et al. (2014), and $z_{t}=0.60^{+0.21}_{-0.12}$ by Yang & Gong (2020). The consistency between the curves from Hubble and Pantheon+ data underscores the robustness of the $f(\mathcal{G})$ model in capturing the expansion dynamics of the Universe.
Figure 5: Graphical representation of the EoS parameter versus redshift using the constrained coefficients from Fig. 1. The thick line represents the behavior of the EoS parameter for the $f(\mathcal{G})$ model, while the dashed line shows the EoS parameter for the $\Lambda$CDM model.
The deceleration parameter is one of the key factors that characterize the behavior of the Universe, determining whether it continuously decelerates, accelerates, or undergoes multiple phases of transition. Similarly, energy sources influence the evolution of the Universe through the EoS parameter, defined as $\omega_{\text{DE}}$ is shown in Fig. 5. By calculating the energy density and pressure of DE, as depicted in Fig. 5, we can observe the variations in the effective EoS of DE relative to the redshift variable. The current EoS values for DE, $\omega_{\text{DE}}(z=0)$, are obtained as $-1.018$, $-0.999$ for the CC, Pantheon+, datasets, respectively. These values indicate phantom behavior (at $z<0$) and a trend towards approximately $-1.32$ at late times. The present values of $\omega_{\text{eff}}$ are $-0.684$, $-0.666$ for the CC and Pantheon+ datasets, respectively. Various cosmological studies have also constrained the EoS parameter, including the Supernovae Cosmology Project $\omega_{\text{DE}}=-1.035^{+0.055}_{-0.059}$ (Amanullah et al., 2010), Planck 2018 $\omega_{\text{DE}}=-1.03\pm 0.03$ (Aghanim et al., 2020), and WAMP+CMB $\omega_{\text{DE}}=-1.079^{+0.090}_{-0.089}$ (Hinshaw et al., 2013).
@@
Fig. 5 shows the effective EoS, $\omega_{\text{eff}}$ as a function of redshift $z$ for the $f(\mathcal{G})$ gravity model. At low redshifts ($z\approx 0$), $\omega_{\text{eff}}$ is close to $-1$, indicating DE dominance and accelerated expansion. As $z$ increases, $\omega_{\text{eff}}$ transitions from negative values to less negative values, reflecting a shift from a deceleration phase in the early Universe to an acceleration-dominated phase in the current epoch. The consistent behavior of $\omega_{\text{DE}}$ across both datasets reinforces the capability of the $f(\mathcal{G})$ model to describe the expansion history of the Universe. This transition aligns with theoretical expectations of the $f(\mathcal{G})$ model, which incorporates higher-order curvature terms to account for cosmic dynamics without a cosmological constant. The observed $\omega_{\text{DE}}$ behavior highlights the effectiveness of the model, supporting its viability as an alternative to the $\Lambda$CDM model.
@@
The statefinder diagnostic proposed by V. Sahni (Sahni et al., 2003) provides a geometric method for discerning different DE models using statefinder parameters.
@@
Fig. 6 illustrates the $r-s$ parameter plot for the $f(\mathcal{G})$ gravity model. The trajectory in the $r-s$ plane highlights the evolutionary track of the expansion of the Universe. The $f(\mathcal{G})$ model passes through the region corresponding to the $\Lambda$CDM model, indicated by the red point. At lower values of $s$, the model aligns with quintessence characteristics, suggesting a dynamical DE component with $\omega>-1$. As $s$ increases, the trajectory moves towards regions associated with Chaplygin gas models, indicating a unified dark matter and DE scenario. The smooth transition observed in the $r-s$ parameter space demonstrates the flexibility of the $f(\mathcal{G})$ model in describing different cosmological behaviors. This capability allows the model to account for various dynamics, from quintessence-like to Chaplygin gas-like, providing a comprehensive description of the expansion of the Universe.
The $\text{Om}(z)$ diagnostic is a simple testing method that depends only on the first-order derivative of the cosmic scale factor. In particular, in DE theories, the $\text{Om}(z)$ parameter is followed as an additional effective diagnostic tool (Sahni et al., 2008, 2003) and is defined as
$\text{Om}(z)=\frac{\frac{H^{2}(z)}{H^{2}_{0}}-1}{(1+z)^{3}-1}.$ (27)
@@
$\text{Om}(z_{1},z_{2})=\text{Om}(z_{1})-\text{Om}(z_{2})\,,$ (28)
Alternatively stated for simplification, when $\text{Om}(z_{1},z_{2})>0$, it indicates quintessence, whereas when $\text{Om}(z_{1},z_{2})<0$, it signifies phantom behavior, where $(z_{1}<z_{2})$. The $\text{Om}(z)$ diagnostic in the $\Lambda$CDM model serves as a null test, as noted in (Sahni et al., 2008), and its sensitivity to the EoS parameter was further explored in subsequent data as seen in (Qi et al., 2018; Zheng et al., 2016; Ding et al., 2015). If $\text{Om}(z)$ remains constant for the redshift, the DE concept would be a cosmological constant. The DE concept will form a cosmological constant if $\text{Om}(z)$ is constant for the redshift. The slope of $\text{Om}(z)$, which is positive for the emerging $\text{Om}(z)$ and denotes phantom phase $(\omega<-1)$ and negative for quintessence region $(\omega>-1)$ also identifies the DE models.
@@
4 Stability Assessment via Dynamical Systems
The techniques of dynamical systems are helpful in examining the overall long-term behavior of a specific cosmological model (Carloni et al., 2005; Amendola et al., 2007; Ivanov & Toporensky, 2012; Carloni et al., 2015; Duchaniya et al., 2023, 2022) and can assist in circumventing the challenge of solving non-linear cosmological equations. These techniques characterize the universal dynamics by studying the nearby long-term behavior of critical points of the system and connecting them to the primary cosmological eras (Paliathanasis, 2017).
@@ 4 Stability Assessment via Dynamical Systems
$X=H^{2}f_{\mathcal{G}}\,,\quad Y=H\dot{f}_{\mathcal{G}}\,,\quad Z=\frac{\dot{H}}{H^{2}}\,,\quad W=-\frac{f}{6H^{2}}\,,\quad V=\frac{\kappa^{2}\rho_{r}}{3H^{2}}\,.$ (29)
alongside the density parameters:
$\Omega_{\text{m}}=\frac{\kappa^{2}\rho_{\text{m}}}{3H^{2}},\quad\Omega_{\text{r}}=V=\frac{\kappa^{2}\rho_{\text{r}}}{3H^{2}},\quad\Omega_{\text{DE}}=\frac{\kappa^{2}\rho_{\text{DE}}}{3H^{2}},$ (30)
@@
Critical Point X Y Z V Existence $\omega_{\text{eff}}$ $q$
$\mathcal{P}_{1}=(x_{1},y_{1},z_{1},v_{1})$ $x_{1}$ $4x_{1}$ $-2$ $1+24x_{1}$
$x_{1}\neq 0$,   $m=\frac{1}{4}$
$\frac{1}{3}$ $1$
$x_{2}\neq 0$,   $m=\frac{1}{4}$
00 $\frac{1}{2}$
$\mathcal{P}_{3}=(x_{3},y_{3},z_{3},v_{4})$ $\frac{-1}{16}$ $\frac{-1}{8}$ $-1$ 00
$2m^{2}-m\neq 0$
$-\frac{1}{3}$ 00
$\mathcal{P}_{4}=(x_{4},y_{4},z_{4},v_{4})$ $\frac{m}{-4+8m}$ 00 00 00
$(-1+2m)(-1+4m)\neq 0$, $m\neq 0$
$-1$ $-1$
$\mathcal{P}_{1}$ 00 $1+24x_{1}$ $-24x_{1}$ Never
$\mathcal{P}_{2}$ $1+20x_{2}$ 00 $-20x_{2}$ Never
$\mathcal{P}_{3}$ 00 00 $1$ Never
$\mathcal{P}_{4}$ 00 00 $1$ Always
Table 4: The density parameters associated with the critical point.
Critical Point Eigenvalues Stability
$\mathcal{P}_{3}$$\left\{-2,-2,-1,4+\frac{2}{-1+2m}\right\}$ Stable for $\frac{1}{4}<m<\frac{1}{2}$
$\mathcal{P}_{4}$$\left\{-4,-3,\frac{3m-6m^{2}-\sqrt{(1-2m)^{2}(25m-4)m}}{2m(2m-1)},\frac{3m-6m^{2}+\sqrt{(1-2m)^{2}(25m-4)m}}{2m(2m-1)}\right\}$ Stable for $\frac{4}{25}\leq m<\frac{1}{4}$
Table 5: Eigenvalues and stability regime.